### v2

@@ 5 Conclusions
This paper presents the weaknesses of Vision Transformers (ViTs) to adversarial attacks in the context of skin disease image classification. According to this study, ViT can perform insanely well on clean images and achieved the most accurate results. In the case of adversarial perturbations, ViT performs poorly and provides less accuracy than traditional CNNs like ResNet-50 and VGG16. The proposed adversarial training method delivers exceptionally good performance using ViT against attacks from adversaries. After using adversarial training, ViT’s accuracy on adversarial images improved a lot from 27.6% to 90.0%. This demonstrates the model’s ability to learn robust representations with a properly regularized model.
@@
The authors declare that they have no known conflict of interests or personal relationships that could have appeared to influence the work reported in this paper.

### v3

@@ 5 Conclusions
This paper presents the weaknesses of Vision Transformers (ViTs) to adversarial attacks in the context of skin disease image classification. According to this study, ViT can perform extremely well on clean images and achieves the most accurate results. However, under adversarial perturbations, ViT performs poorly and provides lower accuracy than traditional CNNs such as ResNet-50 and VGG16. The proposed adversarial training method delivers exceptionally strong performance for ViT against adversarial attacks. After adversarial training, ViT’s accuracy on adversarial images improved significantly from 27.6% to 90.0%. This demonstrates the model’s ability to learn robust representations when trained with properly regularized adversarial objectives.
@@
The authors declare that they have no known conflicts of interest or personal relationships that could have appeared to influence the work reported in this paper.